# CADmium: Fine-Tuning Code Language Models for Text-Driven Sequential CAD Design

### v1

@@ Abstract
Computer-aided design (CAD) is the digital construction of 2D and 3D objects, and is central to a wide range of engineering and manufacturing applications like automobile and aviation. Despite its importance, CAD modeling remains largely a time-intensive, manual task. Recent works have attempted to automate this process with small transformer-based models and handcrafted CAD sequence representations. However, there has been little effort to leverage the potential of large language models (LLMs) for sequential CAD design. In this work, we introduce a new large-scale dataset of more than 170k CAD models annotated with high-quality, human-like descriptions generated with our pipeline based on GPT-4.1. Using this dataset, we fine-tune powerful code-LLMs to generate CAD sequences represented in a JSON-based format from natural language descriptions, demonstrating the viability and effectiveness of this approach for text-conditioned CAD generation. Because simple metrics often fail to reflect the quality of generated objects, we introduce geometric and topological metrics based on sphericity, mean curvature, and Euler characteristic to provide richer structural insights. Our experiments and ablation studies on both synthetic and human-annotated data demonstrate that CADmium is able to automate CAD design, drastically speeding up the design of new objects. The dataset, code, and fine-tuned models are available online  [^1] The code is available at https://github.com/chandar-lab/CADmium, and the models and dataset are available at https://huggingface.co/collections/chandar-lab/cadmium-6866b402be81f39321af98d4. .
**footnotetext: Equal contribution.
@@ 1 Introduction
Recognizing that CAD is a sequential process involving 2D sketches, such as lines and circles, followed by 3D operations like extrusions, recent studies have leveraged transformer architectures and publicly available CAD datasets to automatically generate design steps in an autoregressive manner [1, 2]. Recently, there has been a growing interest in generating 3D objects from natural language descriptions, as this approach would make CAD accessible to a broader, non-expert audience and speed up the creation process even further. Such an interest likely stems from the progress in large pre-trained foundation models that has resulted in tremendous reasoning and generative capabilities in natural language [3] and images [4]. However, unlike these modalities, text-conditioned CAD object generation has not seen the same success as the field still faces two key challenges: (1) the lack of high-quality, human-like textual annotations for CAD samples, and (2) the absence of suitable embeddings for CAD design histories that facilitate the effective use of pre-trained language models.
@@
Current approaches address the first challenge by automating text annotation using sophisticated state-of-the-art annotation pipelines with vision-language and mixture of experts models [2, 5] and the second with vectorized representations of CAD sequences combined with separate embedding layers and prompt tokens to train language models [2, 5]. However, we identify critical shortcomings with both strategies. First, current machine-generated annotations often fail to strike a balance between fluency and geometric specificity, a balance essential for natural human interaction while containing sufficient geometric detail to unambiguously specify the desired 3D object. Second, prior approaches treat CAD generation as a domain-specific task requiring custom embeddings, thereby failing to fully leverage the potential of pre-trained language models and requiring significant computational resources to learn new embeddings from limited data.
@@
To address these limitations, we introduce CADmium [^2] The name CADmium reflects the integration of “CAD” (Computer-Aided Design) and draws inspiration from the metal cadmium, known for its malleability and utility in precision applications., a novel approach that reformulates CAD generation as a purely text-to-text task. Our contributions are as follows. First, we develop a novel annotation pipeline using GPT-4.1 [6] that processes up to 10 multi-view images of 3D objects along with their construction sequences to generate annotations that combine natural language fluency with geometric precision. We apply our pipeline on 176,017 CAD sequences from DeepCAD [1] and release a new dataset with improved expert-quality annotations. Second, we fine-tune Qwen 2.5-Coder-14B [7], a state-of-the-art instruction-tuned code LLM, on our dataset to generate JSON-formatted CAD sequences from natural language prompts, leveraging the capabilities of the pre-trained model without the need for specialized embedding layers. Recent works have shown that fine-tuning LLMs can be effective for generating code used in visual content creation [8, 9], emphasizing their versatility in visual generative applications. We evaluate our approach against state-of-the-art CAD generation models, including Text2CAD [2], using both point-cloud and mesh-based metrics established in prior work. We also validate our models on the Fusion360 Reconstruction dataset [10] using annotations generated by our pipeline and the CADPrompt benchmark [11] that contains 200 objects annotated by human experts. Third, we introduce novel metrics based on sphericity, mean curvature, Euler characteristic, and watertightness to more extensively evaluate the structural and topological characteristics of the generated objects. We highlight the human-like fluency and conciseness of CADmium’s expert-level text annotations relative to Text2CAD, supported by extensive empirical analysis. Furthermore, our fine-tuned LLM achieves competitive performance against the pre-trained Text2CAD model, with improvements across some evaluation metrics as model size increases.
@@ 2 Related Works
The Fusion 360 Gallery [10] was the first dataset of human-designed CAD objects to include their corresponding construction sequences. However, the Fusion 360 Gallery only contains 8,625 samples. As this was not enough for training, DeepCAD [1] introduced a significantly larger dataset of 176,017 CAD models constructed using only sketch and extrude operations parsed from the ABC dataset [12]. Leveraging their large-scale dataset, DeepCAD trained a transformer-based autoencoder within a latent-GAN framework for unconditional CAD generation. SkexGen [13] and SolidGen [14] later improved upon DeepCAD by introducing alternative input representations such as disentangled codebooks and boundary representations. However, these approaches were limited by their unconditional nature, offering no mechanism for users to control the generated objects.
As a result, subsequent works shifted toward conditional generation from point clouds, images, or text inputs [5, 15, 16, 17, 2, 18]. Progress in text-conditioned CAD generation was, however, hindered by the lack of datasets with textual annotations. As a solution, Text2CAD [2] proposed a novel annotation pipeline based on LLMs and vision-language models (VLMs) to process multi-view images and design histories from the DeepCAD dataset and generate annotations ranging from beginner to expert level. Text2CAD produced approximately 660,000 annotations covering all DeepCAD samples and is the only open-source dataset with geometrically precise annotations currently available for training text-conditioned CAD generation models. Nonetheless, beginner-level prompts often lack the details required to unambiguously reconstruct objects, and expert-level prompts are drastically verbose, redundant, or sometimes superfluous. Furthermore, the OmniCAD dataset of machine-generated text annotations for CAD objects, proposed by Xu et al. [5], contains only beginner-level prompts.
In addition to annotating the DeepCAD dataset, Text2CAD introduced a supervised generation pipeline using a frozen BERT encoder [19], adaptive text layers, and a transformer decoder with cross-attention between text and CAD embeddings. CAD-MLLM [5] improved over this framework by incorporating LLMs and conditioning the generations on multiple modalities, namely point clouds, images, and text. CAD-MLLM also introduced a new annotation pipeline and additional CAD samples parsed from the ABC dataset. In parallel, FlexCAD [20] proposed a hierarchical text representation of CAD sequences to make the generation more controllable. With CADmium, we primarily aim to advance state-of-the-art performance in the text-to-CAD task, leaving multimodal integration and enhanced controllability for future work.
@@ 3.1 Datasets and Annotation Pipeline
We address the quality concerns of Text2CAD annotations by re-annotating all 176,017 samples from the DeepCAD library using the same multimodal inputs as Text2CAD. These inputs include the minimal JSON metadata extracted from DeepCAD using Onshape’s FeatureScript [21], which improves human readability by eliminating redundant information, and 10 multi-view images rendered with Blender’s Python module [22] corresponding to the top, bottom, and 8 sides. We leverage the multimodal capabilities of GPT-4.1 to generate concise, natural-sounding yet geometrically precise descriptions. Additionally, we annotate 8,307 CAD objects from the Fusion 360 Reconstruction Dataset [10] not found in the ABC corpus from which DeepCAD is derived, offering a robust measure of generalization. This comprehensive re-annotation effort across both datasets involved processing a total of over 600 million input and output tokens with GPT-4.1. Importantly, only the expert-level annotations in Text2CAD provide enough geometric detail to unambiguously reconstruct objects. While it may be useful to generate plausible objects from incomplete descriptions in some cases, evaluating them is challenging as multiple outputs could satisfy a partial description. As a result, we generate exclusively expert-level prompts and restrict our comparisons to the expert-level descriptions in Text2CAD.
@@
Figure 2: Comparative analysis of CADmium and Text2CAD expert-level annotations. (a) Vocabulary growth as a function of token count demonstrates that Text2CAD has a limited vocabulary compared to CADmium. (b) Human-likeness, clarity, visual faithfulness given image renders, and completeness against the minimal JSON as measured by Gemma-3 12B [23] indicates that CADmium descriptions tend to produce more natural-sounding albeit challenging descriptions. (c-e) Distribution of word counts, unique words, and digit lengths within numerical expressions per annotation shows that CADmium descriptions are more concise and diverse.
Looking first at corpus-level statistics, CADmium’s annotations are drastically more concise, most being between 100 and 200 words, compared to Text2CAD’s longer and often redundant annotations, often exceeding 200 words (Figure  2c). Despite being shorter, CADmium’s annotations achieve a greater lexical variety, with a higher ratio of unique words per annotation (Figure  2d). The vocabulary growth curve also shows that CADmium uses a much broader range of words across annotations, reaching over 18,000 unique words or twice as many as Text2CAD (Figure  2a). This richer vocabulary creates more diverse descriptions and reduces the risk of overfitting to a limited set of words. Additionally, CADmium’s annotations represent numbers in a more natural, human-like way compared to the excessive precision of Text2CAD (Figure  2e). A side-by-side comparison between CADmium and Text2CAD descriptions is provided in Appendix A.
While statistics provide a general feel for the quality of the corpus, they may not tell how individual descriptions compare head-to-head. Therefore, we conducted an LLM-as-a-judge evaluation using Gemma-3 12B [23] on 10,000 annotation pairs (one from CADmium, one from Text2CAD). We looked at four key factors: human-likeness, clarity and readability, visual faithfulness to image renders, and completeness with respect to the minimal JSON description. CADmium was preferred for human-likeness and clarity while achieving comparable visual faithfulness. Text2CAD scored higher on completeness, which aligns with our design choice to prioritize natural-sounding concise annotations over the exhaustive, JSON-like detail found in Text2CAD. We argue that such verbosity, while improving completeness, is less representative of how humans naturally describe CAD models and may artificially simplify the task for language models. Details on the evaluation setup and prompts can be found in Appendix C.
@@ 3.2 Model Selection and Training Details
All models were fine-tuned using AdamW with the default $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. We use a cosine learning rate schedule with 100 warmup steps, reaching a peak learning rate of $2\times 10^{-4}$, and subsequently decaying to a final rate of 00 at the end of the training. To improve computational efficiency, mixed-precision training with fp16 was applied alongside Low-Rank Adaptation (LoRA) [24] across all linear layers, configured with a rank of 64 and an alpha value of 32.
All experiments were conducted on nodes with four NVIDIA A100 80GB GPUs, utilizing Fully Sharded Data Parallel (FSDP) [25] for efficient distributed training, and achieving an effective batch size of 16. The loss function was computed exclusively on the assistant’s generated JSON responses. The models were trained for up to 4 epochs (approximately 29,000 steps), with each complete training run typically lasting between 8 to 12 hours, depending on the specific model size being fine-tuned.
@@ 3.3 Evaluation
where $S$ is the surface of a sphere of volume $V$. Given that a sphere achieves the minimum surface area given a particular volume, $\psi\in(0,1]$. Sphericity, which is explored in foundational work on particle shape characterization [26], and used in various applications for quality assessment of 3D parts, is $1$ for a perfect sphere and decreases for those that are less compact or have more intricate surfaces relative to their volume. The final sphericity discrepancy metric quantifies the difference between the predicted object ($pred$) and the ground truth object ($gt$) as the absolute difference of their normalized ratios:
@@
The second metric we introduce, termed the Discrete Mean Curvature Difference (DMCD), quantifies local dissimilarities in surface geometry. To achieve this, we compute a discrete mean curvature value at each vertex for both the predicted and ground truth meshes. This computation follows the principles for discrete curvature estimation presented by Cohen-Steiner and Morvan [27], leveraging an implementation available in the trimesh library [28]. Specifically, for each vertex, the discrete mean curvature value $\kappa$ is calculated as the sum of the signed dihedral angles of all mesh edges contained within a sphere of a defined radius $r$ centered at that vertex. This per-vertex value $\kappa$ is inherently dimensionless. The curvature at each vertex is then averaged, resulting in $\bar{\kappa}$. We compute the absolute difference between the average $\kappa$ values of the predicted and ground truth meshes:
@@
We also introduce the Exact Euler Characteristic Match (EECM) metric. The Euler characteristic, denoted as $\chi$, is a well-established topological invariant that describes a shape’s structure regardless of how it is bent or deformed. For a 3D manifold mesh, it is commonly computed using the formula $\chi=V-E+F$, where $V$ is the number of vertices, $E$ the number of edges, and $F$ the number of faces [29]. The EECM is a binary metric defined as
@@
A fundamental quality indicator for generated 3D shapes, especially those intended for physical simulation, manufacturing, or volumetric analysis, is whether they constitute a Watertight Mesh. A watertight mesh is a closed, 2-manifold polygonal surface that clearly separates an interior volume from an exterior space, meaning it has no holes, boundary edges, or non-manifold configurations that would prevent the unambiguous definition of its volume [30]. We report the percentage of generated CAD models that are watertight as a measure of their geometric validity and usability. In our evaluation pipeline, we compute SD, EECM, and DMCD only if both the predicted and ground truth meshes are watertight.

### v2

@@ Abstract
Computer-aided design (CAD) is the digital construction of 2D and 3D objects, and is central to a wide range of engineering and manufacturing applications like automobile and aviation. Despite its importance, CAD modeling remains largely a time-intensive, manual task. Recent works have attempted to automate this process with small transformer-based models and handcrafted CAD sequence representations. However, there has been little effort to leverage the potential of large language models (LLMs) for sequential CAD design. In this work, we introduce a new large-scale dataset of more than 170k CAD models annotated with high-quality, human-like descriptions generated with our pipeline based on GPT-4.1. Using this dataset, we fine-tune powerful code-LLMs to generate CAD sequences represented in a JSON-based format from natural language descriptions, demonstrating the viability and effectiveness of this approach for text-conditioned CAD generation. Because simple metrics often fail to reflect the quality of generated objects, we introduce geometric and topological metrics based on sphericity, mean curvature, and Euler characteristic to provide richer structural insights. Our experiments and ablation studies on both synthetic and human-annotated data demonstrate that CADmium is able to automate CAD design, drastically speeding up the design of new objects. The dataset, code, and fine-tuned models are available online.
Data and Model: huggingface.co/collections/chandar-lab/cadmium-6866b402be81f39321af98d4
Code Repository: github.com/chandar-lab/cadmium
@@ 1 Introduction
Recognizing that CAD is a sequential process involving 2D sketches, such as lines and circles, followed by 3D operations like extrusions, recent studies have leveraged transformer architectures and publicly available CAD datasets to automatically generate design steps in an autoregressive manner (Wu et al., 2021; Khan et al., 2024). Recently, there has been a growing interest in generating 3D objects from natural language descriptions, as this approach would make CAD accessible to a broader, non-expert audience and speed up the creation process even further. Such an interest likely stems from the progress in large pre-trained foundation models that has resulted in tremendous reasoning and generative capabilities in natural language (Brown et al., 2020) and images (Esser et al., 2024). However, unlike these modalities, text-conditioned CAD object generation has not seen the same success as the field still faces two key challenges: (1) the lack of high-quality, human-like textual annotations for CAD samples, and (2) the absence of suitable embeddings for CAD design histories that facilitate the effective use of pre-trained language models.
@@
Current approaches address the first challenge by automating text annotation using sophisticated state-of-the-art annotation pipelines with vision-language and mixture of experts models (Khan et al., 2024; Xu et al., 2024) and the second with vectorized representations of CAD sequences combined with separate embedding layers and prompt tokens to train language models (Khan et al., 2024; Xu et al., 2024). However, we identify critical shortcomings with both strategies. First, current machine-generated annotations often fail to strike a balance between fluency and geometric specificity, a balance essential for natural human interaction while containing sufficient geometric detail to unambiguously specify the desired 3D object. Second, prior approaches treat CAD generation as a domain-specific task requiring custom embeddings, thereby failing to fully leverage the potential of pre-trained language models and requiring significant computational resources to learn new embeddings from limited data.
@@
To address these limitations, we introduce CADmium [^1] The name CADmium reflects the integration of “CAD” (Computer-Aided Design) and draws inspiration from the metal cadmium, known for its malleability and utility in precision applications., a novel approach that reformulates CAD generation as a purely text-to-text task. Our contributions are as follows. First, we develop a novel annotation pipeline using GPT-4.1 (OpenAI, 2025) that processes up to 10 multi-view images of 3D objects along with their construction sequences to generate annotations that combine natural language fluency with geometric precision. We apply our pipeline on 176,017 CAD sequences from DeepCAD (Wu et al., 2021) and release a new dataset with improved expert-quality annotations. Second, we fine-tune Qwen 2.5-Coder-14B (Yang et al., 2024), a state-of-the-art instruction-tuned code LLM, on our dataset to generate JSON-formatted CAD sequences from natural language prompts, leveraging the capabilities of the pre-trained model without the need for specialized embedding layers. Recent works have shown that fine-tuning LLMs can be effective for generating code used in visual content creation (Rodriguez et al., 2024, 2025), emphasizing their versatility in visual generative applications. We evaluate our approach against state-of-the-art CAD generation models, including Text2CAD (Khan et al., 2024), using both point-cloud and mesh-based metrics established in prior work. We also validate our models on the Fusion360 Reconstruction dataset (Willis et al., 2021) using annotations generated by our pipeline and the CADPrompt benchmark (Alrashedy et al., 2024) that contains 200 objects annotated by human experts. Third, we introduce novel metrics based on sphericity, mean curvature, Euler characteristic, and watertightness to more extensively evaluate the structural and topological characteristics of the generated objects. We highlight the human-like fluency and conciseness of CADmium’s expert-level text annotations relative to Text2CAD, supported by extensive empirical analysis. Furthermore, our fine-tuned LLM achieves competitive performance against the pre-trained Text2CAD model, with improvements across some evaluation metrics as model size increases.
@@ 2 Related Works
The Fusion 360 Gallery (Willis et al., 2021) was the first dataset of human-designed CAD objects to include their corresponding construction sequences. However, the Fusion 360 Gallery only contains 8,625 samples. As this was not enough for training, DeepCAD (Wu et al., 2021) introduced a significantly larger dataset of 176,017 CAD models constructed using only sketch and extrude operations parsed from the ABC dataset (Koch et al., 2019). Leveraging their large-scale dataset, DeepCAD trained a transformer-based autoencoder within a latent-GAN framework for unconditional CAD generation. SkexGen (Xu et al., 2022) and SolidGen (Jayaraman et al., 2022) later improved upon DeepCAD by introducing alternative input representations such as disentangled codebooks and boundary representations. However, these approaches were limited by their unconditional nature, offering no mechanism for users to control the generated objects.
As a result, subsequent works shifted toward conditional generation from point clouds, images, or text inputs (Xu et al., 2024; You et al., 2024; Alam & Ahmed, 2024; Uy et al., 2022; Khan et al., 2024; Dupont et al., 2024). Progress in text-conditioned CAD generation was, however, hindered by the lack of datasets with textual annotations. As a solution, Text2CAD (Khan et al., 2024) proposed a novel annotation pipeline based on LLMs and vision-language models (VLMs) to process multi-view images and design histories from the DeepCAD dataset and generate annotations ranging from beginner to expert level. Text2CAD produced approximately 660,000 annotations covering all DeepCAD samples and is the only open-source dataset with geometrically precise annotations currently available for training text-conditioned CAD generation models. Nonetheless, beginner-level prompts often lack the details required to unambiguously reconstruct objects, and expert-level prompts are drastically verbose, redundant, or sometimes superfluous. Furthermore, the OmniCAD dataset of machine-generated text annotations for CAD objects, proposed by Xu et al. (2024), contains only beginner-level prompts.
In addition to annotating the DeepCAD dataset, Text2CAD introduced a supervised generation pipeline using a frozen BERT encoder (Devlin et al., 2019), adaptive text layers, and a transformer decoder with cross-attention between text and CAD embeddings. CAD-MLLM (Xu et al., 2024) improved over this framework by incorporating LLMs and conditioning the generations on multiple modalities, namely point clouds, images, and text. CAD-MLLM also introduced a new annotation pipeline and additional CAD samples parsed from the ABC dataset. In parallel, FlexCAD (Zhang et al., 2024) proposed a hierarchical text representation of CAD sequences to make the generation more controllable. With CADmium, we primarily aim to advance state-of-the-art performance in the text-to-CAD task, leaving multimodal integration and enhanced controllability for future work.
@@ 3.1 Datasets and Annotation Pipeline
We address the quality concerns of Text2CAD annotations by re-annotating all 176,017 samples from the DeepCAD library using the same multimodal inputs as Text2CAD. These inputs include the minimal JSON metadata extracted from DeepCAD using Onshape’s FeatureScript (ons, ), which improves human readability by eliminating redundant information, and 10 multi-view images rendered with Blender’s Python module (Blender Online Community, 2025) corresponding to the top, bottom, and 8 sides. We leverage the multimodal capabilities of GPT-4.1 to generate concise, natural-sounding yet geometrically precise descriptions. Additionally, we annotate 8,307 CAD objects from the Fusion 360 Reconstruction Dataset (Willis et al., 2021) not found in the ABC corpus from which DeepCAD is derived, offering a robust measure of generalization. This comprehensive re-annotation effort across both datasets involved processing a total of over 600 million input and output tokens with GPT-4.1. Importantly, only the expert-level annotations in Text2CAD provide enough geometric detail to unambiguously reconstruct objects. While it may be useful to generate plausible objects from incomplete descriptions in some cases, evaluating them is challenging as multiple outputs could satisfy a partial description. As a result, we generate exclusively expert-level prompts and restrict our comparisons to the expert-level descriptions in Text2CAD.
@@
Figure 2: Comparative analysis of CADmium and Text2CAD expert-level annotations. (a) Vocabulary growth as a function of token count demonstrates that Text2CAD has a limited vocabulary compared to CADmium. (b) Human-likeness, clarity, visual faithfulness given image renders, and completeness against the minimal JSON as measured by Gemma-3 12B (Team et al., 2025) indicates that CADmium descriptions tend to produce more natural-sounding albeit challenging descriptions. (c-e) Distribution of word counts, unique words, and digit lengths within numerical expressions per annotation shows that CADmium descriptions are more concise and diverse.
Looking first at corpus-level statistics, CADmium’s annotations are drastically more concise, most being between 100 and 200 words, compared to Text2CAD’s longer and often redundant annotations, often exceeding 200 words (Figure 2c). Despite being shorter, CADmium’s annotations achieve a greater lexical variety, with a higher ratio of unique words per annotation (Figure 2d). The vocabulary growth curve also shows that CADmium uses a much broader range of words across annotations, reaching over 18,000 unique words or twice as many as Text2CAD (Figure 2a). This richer vocabulary creates more diverse descriptions and reduces the risk of overfitting to a limited set of words. Additionally, CADmium’s annotations represent numbers in a more natural, human-like way compared to the excessive precision of Text2CAD (Figure 2e). A side-by-side comparison between CADmium and Text2CAD descriptions is provided in Appendix A.
While statistics provide a general feel for the quality of the corpus, they may not tell how individual descriptions compare head-to-head. Therefore, we conducted an LLM-as-a-judge evaluation using Gemma-3 12B (Team et al., 2025) on 10,000 annotation pairs (one from CADmium, one from Text2CAD). We looked at four key factors: human-likeness, clarity and readability, visual faithfulness to image renders, and completeness with respect to the minimal JSON description. CADmium was preferred for human-likeness and clarity while achieving comparable visual faithfulness. Text2CAD scored higher on completeness, which aligns with our design choice to prioritize natural-sounding concise annotations over the exhaustive, JSON-like detail found in Text2CAD. We argue that such verbosity, while improving completeness, is less representative of how humans naturally describe CAD models and may artificially simplify the task for language models. Details on the evaluation setup and prompts can be found in Appendix C.
@@ 3.2 Model Selection and Training Details
All models were fine-tuned using AdamW with the default $\beta_{1}=0.9$, $\beta_{2}=0.999$, and $\epsilon=10^{-8}$. We use a cosine learning rate schedule with 100 warmup steps, reaching a peak learning rate of $2\times 10^{-4}$, and subsequently decaying to a final rate of 0 at the end of the training. To improve computational efficiency, mixed-precision training with fp16 was applied alongside Low-Rank Adaptation (LoRA) (Hu et al., 2022) across all linear layers, configured with a rank of 64 and an alpha value of 32.
All experiments were conducted on nodes with four NVIDIA A100 80GB GPUs, utilizing Fully Sharded Data Parallel (FSDP) (Zhao et al., 2023) for efficient distributed training, and achieving an effective batch size of 16. The loss function was computed exclusively on the assistant’s generated JSON responses. The models were trained for up to 4 epochs (approximately 29,000 steps), with each complete training run typically lasting between 8 to 12 hours, depending on the specific model size being fine-tuned.
@@ 3.3 Evaluation
where $S$ is the surface of a sphere of volume $V$. Given that a sphere achieves the minimum surface area given a particular volume, $\Psi\in(0,1]$. Sphericity, which is explored in foundational work on particle shape characterization (Wadell, 1932), and used in various applications for quality assessment of 3D parts, is $1$ for a perfect sphere and decreases for those that are less compact or have more intricate surfaces relative to their volume. The final sphericity discrepancy metric quantifies the difference between the predicted object (predpred) and the ground truth object ($gt$) as the absolute difference of their normalized ratios:
@@
The second metric we introduce, termed the Discrete Mean Curvature Difference (DMCD), quantifies local dissimilarities in surface geometry. To achieve this, we compute a discrete mean curvature value at each vertex for both the predicted and ground truth meshes. This computation follows the principles for discrete curvature estimation presented by Cohen-Steiner & Morvan (2003), leveraging an implementation available in the trimesh library (Dawson-Haggerty et al., ). Specifically, for each vertex, the discrete mean curvature value $\kappa$ is calculated as the sum of the signed dihedral angles of all mesh edges contained within a sphere of a defined radius $r$ centered at that vertex. This per-vertex value $\kappa$ is inherently dimensionless. The curvature at each vertex is then averaged, resulting in $\bar{\kappa}$. We compute the absolute difference between the average $\kappa$ values of the predicted and ground truth meshes:
@@
We also introduce the Exact Euler Characteristic Match (EECM) metric. The Euler characteristic, denoted as $\chi$, is a well-established topological invariant that describes a shape’s structure regardless of how it is bent or deformed. For a 3D manifold mesh, it is commonly computed using the formula $\chi=V-E+F$, where $V$ is the number of vertices, $E$ the number of edges, and $F$ the number of faces (Richeson, 2012). The EECM is a binary metric defined as
@@
A fundamental quality indicator for generated 3D shapes, especially those intended for physical simulation, manufacturing, or volumetric analysis, is whether they constitute a Watertight Mesh. A watertight mesh is a closed, 2-manifold polygonal surface that clearly separates an interior volume from an exterior space, meaning it has no holes, boundary edges, or non-manifold configurations that would prevent the unambiguous definition of its volume (Botsch et al., 2010). We report the percentage of generated CAD models that are watertight as a measure of their geometric validity and usability. In our evaluation pipeline, we compute SD, EECM, and DMCD only if both the predicted and ground truth meshes are watertight.